OpenStudy (anonymous):

A rectangle has a length of x + 2, a width of 7, and a perimeter of 32. The value of x is

OpenStudy (anonymous):

Perimeter=2(l+b)

OpenStudy (anonymous):

32=2(x+2+7)

OpenStudy (anonymous):

16=x+9 ---->x=16-9=7

OpenStudy (anonymous):

got it?